# SSTP: Efficient Sample Selection for Trajectory Prediction

### v2

@@ Abstract
Trajectory prediction is a core task in autonomous driving. However, training advanced trajectory prediction models on large-scale datasets is both time-consuming and computationally expensive. In addition, the imbalanced distribution of driving scenarios often biases models toward data-rich cases, limiting performance in safety-critical, data-scarce conditions. To address these challenges, we propose the Sample Selection for Trajectory Prediction (SSTP) framework, which constructs a compact yet balanced dataset for trajectory prediction. SSTP consists of two main stages (1) Extraction, in which a pretrained trajectory prediction model computes gradient vectors for each sample to capture their influence on parameter updates; and (2) Selection, where a submodular function is applied to greedily choose a representative subset that covers diverse driving scenarios. This approach significantly reduces the dataset size and mitigates scenario imbalance, without sacrificing prediction accuracy and even improving in high-density cases. We evaluate our proposed SSTP on the Argoverse 1 and Argoverse 2 benchmarks using a wide range of recent state-of-the-art models. Our experiments demonstrate that SSTP achieves comparable performance to full-dataset training using only half the data while delivering substantial improvements in high-density traffic scenes and significantly reducing training time. Importantly, SSTP exhibits strong generalization and robustness, and the selected subset is model-agnostic, offering a broadly applicable solution. The code is avaiable at https://github.com/RuiningYang/SSTP.
@@ 1 Introduction
Figure 1: Top: Comparison of model performance across varying scene densities. Superior enhancement on HiVT-64 [61] in high-density scenarios training with SSTP-selected 50% data from Argoverse 1 [5]. Bottom: Using the SSTP-selected subset results in a significant overall reduction in training time compared to using the full dataset, especially for complex models like HPNet [42].
@@
Trajectory prediction aims to predict the future locations of agents conditioned on their past observations, which plays a key role in the domain of autonomous driving. This task is essential yet challenging due to the complex uncertain driving situations. With rapid developments in deep learning, various methods [61, 14, 4, 16, 52, 57, 30] have been proposed with promising trajectory prediction performance. Meanwhile, more large-scale realistic datasets [1, 49, 5, 47, 40] have been released by research institute and self-driving companies, which further push the boundary of this task. However, one common issue is that training these data-driven methods requires enormous computational resources and is time-consuming due to the large scale of the datasets. For example, the recent state-of-the-art method MTR [30] has over 66 million model parameters, and the Waymo Open Motion Dataset (WOMD) [40] has over 2.2 million trajectory samples. Training the complete model on this dataset requires a substantial amount of GPU hours, posing a significant computational burden. A similar situation arises in multiple methods as well as across various datasets, which is shown in Table 1. This phenomenon raises two following questions:
Q1: Do we need large-scale datasets to train the model?
@@
Q2: How can we effectively reduce the training data volume without significantly compromising model accuracy?
@@
Prompted by these questions, we take a deep dive into recent trajectory prediction benchmarks and reveal one finding the imbalance in the dataset. From a conventional machine learning perspective, standard tasks typically evaluate model performance using the average accuracy across all samples. However, in autonomous driving trajectory prediction, this metric inherently biases models toward driving scenarios with abundant data, leading to suboptimal performance in data-scarce scenarios. For safe driving, an ideal trajectory predictor should exhibit robust performance across both data-rich and data-scarce scenarios. However, due to the imbalanced distribution of driving scenarios, existing models struggle to maintain consistent performance across diverse scenarios, as illustrated in Figure 1.
Model Dataset Samples Training Time
HiVT [61] Argoverse 1 [5] 330k 43 hours
HPNet [42] Interaction [49] 62k 51 hours
HPNet [42] Argoverse 1 [5] 330k 256 hours
QCNet [62] Argoverse 2 [47] 250k 144 hours
DeMo [50] Argoverse 2 [47] 250k 72 hours
Table 1: Statistics on the training time of trajectory prediction models and the size of the datasets used. All training time data are derived from the original experimental setup and hardware specified by the model authors.
To address this issue, we introduce the Sample Selection for Trajectory Prediction (SSTP) framework, the first framework designed to construct a compact and balanced dataset for trajectory prediction. SSTP consists of two main stages: extraction and selection. In the extraction stage, a baseline model is pretrained on the full dataset for a few epochs to establish an initial feature representation. At the same time, the dataset is partitioned according to the number of agents per sample, and gradient vectors are computed for each sample to capture its influence on parameter updates. In the selection stage, we adopt a greedy strategy with a submodular function to compute scores and iteratively select representative samples. This approach builds a target subset that covers diverse driving scenarios while preserving the most distinctive features of each group. As shown in Figure 1, SSTP uniquely reduces data volume while maintaining, or even improving, model performance, particularly in high-density scenarios. Furthermore, it effectively balances the data distribution across various driving scenarios, making it well-suited for training a wide range of trajectory prediction models.
@@
We evaluate our proposed method on the Argoverse 1 [5] and Argoverse 2 [47] datasets with multiple baseline methods. Empirical results demonstrate that SSTP successfully constructs a compact and well-balanced dataset. In high-density scenarios, training on the selected subset significantly outperforms using the full dataset, demonstrating the effectiveness of our method. Moreover, our work offers a resource-efficient dataset that maintains balanced performance across various driving scenarios, making it well-suited for training state-of-the-art trajectory prediction models. The main contributions are summarized as follows:
  • We introduce SSTP, a novel and effective framework for constructing a compact, balanced dataset for trajectory prediction, which boosts model performance across diverse driving scenarios.
  • SSTP significantly reduces training time and computational cost with a one-time selection process, making it efficient and adaptable to various baseline models.
  • Extensive experiments show that with only 50% of the original data, SSTP achieves comparable, or even better, performance across scenarios compared to full-dataset.
2 Related Work
2.1 Trajectory Prediction
Trajectory prediction infers an agent’s future motion from its historical observations. In recent years, research has increasingly concentrated on capturing complex multi-agent interactions, driving advances in various predictive methods [43, 15, 48, 25, 30, 61, 62, 42, 50, 23]. Furthermore, novel approaches including pretraining [7, 9, 21], historical prediction structure design [32, 42], GPT-style next-token prediction [34, 39], and post-processing optimization [60, 10] have significantly enhanced model performance, demonstrating strong results across various datasets. However, most state-of-the-art methods rely on training with large-scale datasets [1, 5, 40, 47, 13], leading to significant computational costs. In contrast, we propose a pioneering data selection strategy in the trajectory prediction domain that constructs a compact, balanced, yet highly representative dataset, significantly reducing training time while preserving model performance.
2.2 Long-Tail in Trajectory Prediction Dataset
The performance of the trajectory prediction models is evaluated based on the overall average. While they excel on benchmarks, these models often struggle with challengingscenarios [27, 46, 35] due to the long-tail data distribution, where common cases dominate and complex or rare situations are underrepresented [6]. Recent studies [46, 51, 58, 22] have started addressing the long-tail problem in trajectory prediction, primarily by leveraging contrastive learning to enhance feature representations. However, these methods mainly focus on optimizing feature-level learning while overlooking the distribution of data at the scenario level and the importance of individual samples. In contrast, our proposed method assesses the contribution of each sample and applies a refined selection strategy to build a balanced and compact dataset. Experimental results show that it significantly boosts performance in complex scenarios.
2.3 Training Sample Selection
Deep neural networks, especially Transformer-based models, depend on large-scale datasets and incur high computational costs. To reduce these costs and shorten training time, various methods focused on improving data efficiency have been proposed, including frequent parameter updates [37], fewer iterations [41], and dynamic learning rate adjustments [20, 12]. To directly reduce data volume, dataset condensation compresses raw data into compact synthetic samples [45, 55, 53, 19, 44, 54, 3]. Another widely studied approach is coreset selection, which constructs a weighted subset that closely approximates the statistical distribution of the original dataset [17, 11, 28, 29]. However, those approaches on dataset condensation [31, 26, 59, 38, 56] and coreset selection [18, 33] has been predominantly focused on image classification. In contrast, we propose a sample selection strategy based on submodular functions for the trajectory prediction domain. As a pioneering method, our approach significantly reduces the training data required while maintaining model performance.
@@ 3 Method
3.1 Problem Formulation
Consider a training set $\mathcal{D}=\{S_{j}\}_{j=1}^{M}$ consists of total $M$ driving samples. Each sample can be described by a triple $S=(X,Y,\mathcal{O})$, where $X$ denotes the observed trajectories, $Y$ represents the corresponding future trajectories for all the agents present, and $\mathcal{O}$ contains driving context information (e.g., maps). Trajectory prediction aims to estimate the future trajectory $Y$ conditioned on $X$ and $\mathcal{O}$. In this paper, our objective is to identify a smaller subset $\mathcal{C}\subseteq\mathcal{D}$ such that training a model in $\mathcal{C}$ stays within a limited budget $\mathcal{B}$. We define the budget as a fraction $\alpha$ of the full dataset size, i.e., $\mathcal{B}=\alpha|\mathcal{D}|$. At the same time, the model trained on $\mathcal{C}$ can achieve comparable trajectory prediction performance to that of the model trained on the full dataset $\mathcal{D}$.
Data Partitioning. As pointed out in Section 1, the data imbalance situation often occurs in commonly used datasets, specifically in large-scale ones. We can observe from Figure 2 that low-density scenes with few agents constitute a large portion of the dataset, while high-density scenes are underrepresented. This imbalance biases model training, making it more inclined to learn patterns from simpler scenarios while struggling to generalize to high-density scenarios. According to this observation, we first compute a density level $\rho(S_{j})$ for each sample $S_{j}\in\mathcal{D}$ based on the number of agents present. Then, we use a fixed interval $\tau$ to partition the dataset into $K$ disjoint subsets based on these density values. The partitioning can be defined as follows:
[…]
where $\rho_{min}$ is the minimum density level in $\mathcal{D}$.
@@ 3.1 Problem Formulation
2:Initialize: $\mathcal{B}\leftarrow\alpha|\mathcal{D}|$; $\triangleright$ Set budget
@@
6:     $n_{k}\leftarrow\texttt{DynamicBudget}(\mathcal{B},k)$;
@@
15:     $\mathcal{B}\leftarrow\mathcal{B}-|\mathcal{C}_{k}|$; $\triangleright$ Update remaining budget
@@
18:function DynamicBudget($\mathcal{B},k$):
19:     if $|\mathcal{D}_{k}|\leq\lfloor\mathcal{B}/{k}\rfloor$ then:
@@
22:         $n_{k}\leftarrow\lfloor\mathcal{B}/{k}\rfloor$;
@@
where $\mathcal{L}_{reg}$ measures the L2 norm difference between the best mode prediction and the ground truth trajectory, and $\mathcal{L}_{cls}$ aligns the mode predicted probabilities $\pi$ with the best mode.
By backpropagating the loss $\mathcal{L}$, we can obtain the gradient with respect to the model output $\hat{\mathbf{Y}}$ as follows:
@@
Sample Selection. Given total $K$ disjoint subsets and the corresponding gradient feature vectors $\mathcal{G}_{k}$, we present a submodular gain function $P(\cdot)$ to evaluate the contribution of each sample. And we utilize a greedy algorithm to iteratively construct the target dataset $\mathcal{C}$ via optimizing the submodular gain function to ensure the selected target dataset optimally represents the entire dataset.
For each subset $\mathcal{D}_{k}$ with $|\mathcal{D}_{k}|>n_{k}$, where $n_{k}$ is the dynamic budget determined by the DynamicBudget function (see Algorithm 1), we initialize an empty set $\mathcal{C}_{k}$ to store the selected samples. For these subsets, let $\mathcal{G}_{k}$ denote the gradient feature vectors computed for the samples in $\mathcal{D}_{k}$. To evaluate the contribution of each sample, we define a submodular gain function based on a cosine similarity kernel. Specifically, for any trajectory sample $S_{j}\in\mathcal{D}_{k}$, the gain function is defined as follows:
@@
where we use a cosine similarity kernel to measure the similarity between the sample $S_{j}$ and other samples. We then apply a greedy optimization strategy, iteratively selecting the sample follows:
@@
At each iteration, the selected sample $S^{*}$ is added to $\mathcal{C}_{k}$. This process continues until the number of selected samples reaches the budget $n_{k}$ for subset $\mathcal{D}_{k}$. Note that we process the subsets starting with those having higher density levels, as these subsets tend to contain relatively fewer samples and are underrepresented. For subsets $\mathcal{D}_{k}$ where $|\mathcal{D}_{k}|\leq n_{k}$, we directly set $\mathcal{C}_{k}=\mathcal{D}_{k}$.
Finally, we can yield the target dataset $\mathcal{C}=\bigcup_{k=1}^{K}\mathcal{C}_{k}$. By incorporating gradient-based similarity into the submodular selection algorithm, our method ensures that the chosen subset maximally covers the gradient space while maintaining diversity, resulting in a smaller dataset that is both representative and informative, which ultimately leads to improved generalization in trajectory prediction models.

### v3

@@ Abstract
Trajectory prediction is a core task in autonomous driving. However, training advanced trajectory prediction models on existing large-scale datasets is both time-consuming and computationally expensive. More critically, these datasets are highly imbalanced in scenario density, with normal driving scenes (low-moderate traffic) overwhelmingly dominating the datasets, while high-density and safety-critical cases are underrepresented. As a result, models tend to overfit low/moderate-density scenarios and perform poorly in high-density scenarios. To address these challenges, we propose the SSTP framework, which constructs a compact yet density-balanced dataset tailored to trajectory prediction. SSTP consists of two main stages: (1) Extraction, where a baseline model is pretrained for a few epochs to obtain stable gradient estimates, and the dataset is partitioned by scenario density. (2) Selection, where gradient-based scores and a submodular objective select representative samples within each density category, while biased sampling emphasizes rare high-density interactions to avoid dominance by low-density cases. This approach significantly reduces the dataset size and mitigates scenario imbalance, without sacrificing prediction accuracy. Experiments on the Argoverse 1 and Argoverse 2 datasets with recent state-of-the-art models show that SSTP achieves comparable performance to full-dataset training using only half the data while delivering substantial improvements in high-density traffic scenes and significantly reducing training time. Robust trajectory prediction depends not only on data scale but also on balancing scene density to ensure reliable performance under complex multi agent interactions.
@@ 1 Introduction
Trajectory prediction aims to predict the future locations of agents conditioned on their past observations, which plays a key role in the domain of autonomous driving. This task is essential yet challenging due to the complex uncertain driving situations. With rapid developments in deep learning, various methods (Zhou et al., 2022c; Feng et al., 2023; Chai et al., 2019; Gu et al., 2021; Zhang et al., 2021; Zhao et al., 2021; Ngiam et al., 2021) have been proposed with promising trajectory prediction performance. Meanwhile, more large-scale realistic datasets (Caesar et al., 2020; Zhan et al., 2019; Chang et al., 2019; Wilson et al., 2023; Sun et al., 2020) have been released by research institutes and self-driving companies, which further push the boundary of this task.
However, one common issue is that training these data-driven methods requires enormous computational resources and is time-consuming due to the large scale of the datasets. For example, the recent state-of-the-art method MTR (Ngiam et al., 2021) has over 66 million model parameters, and the Waymo Open Motion Dataset (WOMD) (Sun et al., 2020) has over 2.2 million trajectory samples. Training the complete model on this dataset requires a substantial amount of GPU hours, posing a significant computational burden.
@@
Q1: To what extent does training on massive trajectory datasets improve performance, considering their high computational cost? Q2: How can we effectively reduce the training data volume without significantly compromising model accuracy?
@@
Prompted by these questions, we take a deep dive into recent trajectory prediction benchmarks and reveal one finding: the imbalance in the dataset. Specifically, the majority of scenarios involve only a limited number of interacting agents, whereas high-density interaction scenarios are significantly underrepresented. This imbalance is evident in both Argoverse 1 (Chang et al., 2019) and Argoverse 2 (Wilson et al., 2023) (Figure 2), where most cases are low-density. Yet such high-density cases are precisely the most safety-critical for autonomous driving, since they involve complex multi-agent interactions where small prediction errors may directly compromise driving safety. Similar long-tail phenomena have been observed in the literature (Makansi et al., 2021; Wang et al., 2023; Pourkeshavarz et al., 2023; Chen et al., 2024). From a conventional machine learning perspective, standard tasks typically evaluate model performance using the average accuracy across all samples. However, in autonomous driving trajectory prediction, this metric inherently biases models toward driving scenarios with abundant data, leading to suboptimal performance in data-scarce scenarios. For safe driving, an ideal trajectory predictor should exhibit robust performance across both data-rich and data-scarce scenarios. However, due to the imbalanced distribution of driving scenarios, state-of-the-art baselines such as HiVT (Zhou et al., 2022c) and HPNet (Tang et al., 2024) struggle to maintain consistent performance across diverse densities, as illustrated in Figure 1.
To address this issue, we introduce the Sample Selection for Trajectory Prediction (SSTP) framework, the first framework designed to construct a compact and balanced dataset for trajectory prediction. SSTP consists of two main stages: extraction and selection. In the extraction stage, a baseline model is pretrained on the full dataset for a few epochs to obtain stable gradient estimates, and the dataset is partitioned by scenario density, measured by the number of agents. This design mitigates the drawbacks of traditional trajectory prediction training: in rare high-density scenarios, the scarcity of training data leads to systematic under-training. For Transformer-based architectures, this issue is particularly pronounced. High-density scenarios require a model to accurately grasp the complex agent–agent dependencies, yet are significantly underrepresented in the training dataset. Consequently, the model does not have sufficient chances to learn such dependencies, leading to poor performance at inference. Moreover, in standard Transformer self-attention, parameters are shared across samples and rely on broad coverage for generalization. When high-density scenarios are scarce, updates are dominated by low-density scenarios, further reinforcing systematic under-training. In the selection stage, for each density category, we calculate gradient-based influence scores for every sample, then apply a submodular objective that selects a subset that captures the most representative cases while reducing redundancy. Across partitions, we employ biased sampling to explicitly upweight rare but critical high-density scenarios, preventing dominance by the majority of low-density scenarios. As shown in Figure 1, SSTP reduces data volume while maintaining performance, and achieves even better results in high-density scenarios.
Figure 1: Comparison of minADE across different scenario densities for HiVT and HPNet on Argoverse 1. Models are trained with the full dataset, 50% randomly selected data, 50% cluster-based selected data, and 50% SSTP (ours). SSTP consistently achieves lower errors in high-density scenarios.
@@
We evaluate our proposed method on the Argoverse 1 (Chang et al., 2019) and Argoverse 2 (Wilson et al., 2023) datasets with multiple baseline methods. Empirical results demonstrate that SSTP successfully constructs a compact and well-balanced dataset. In high-density scenarios, training on the selected subset significantly outperforms using the full dataset, demonstrating the effectiveness of our method. Moreover, our work offers a resource-efficient dataset that maintains balanced performance across various driving scenarios, making it well-suited for training state-of-the-art trajectory prediction models. The main contributions are summarized as follows:
  • We reveal that modern trajectory prediction datasets are heavily skewed, with scene density and other long-tailed factors biasing models toward majority regimes and limiting performance in safety-critical, rare scenarios.
  • We introduce SSTP, a framework that partitions data by scene density and applies submodular selection with gradient-based influence scores, combined with biased sampling to emphasize underrepresented yet safety-critical cases.
  • By selecting only 50% of the data, SSTP achieves performance comparable to or better than full-data training, substantially reduces computational cost, reduces error in high-density interaction scenes, and transfers robustly across diverse model architectures.
Figure 2: Visualization of different-density scenarios in trajectory prediction datasets. Left shows examples from Argoverse 1, and right shows examples from Argoverse 2.
@@ 2 Method
Denote the training set as $\mathcal{D}=\{S_{j}\}_{j=1}^{M}$. Each sample is represented as a triple $S=(X,Y,\mathcal{O})$, where $X$ and $Y$ denote the observed trajectories and future trajectories of all agents, and $\mathcal{O}$ the driving context information (e.g., maps). The goal is to estimate $Y$ conditioned on $X$ and $\mathcal{O}$. We aim to construct a compact subset $\mathcal{C}\subseteq\mathcal{D}$ of target size $B$.
@@
2:Initialize: $B\leftarrow\alpha|\mathcal{D}|$; $\triangleright$ Set target size
@@
6:  $n_{k}\leftarrow\texttt{DynamicSelect}(B,k)$;
@@
15:  $B\leftarrow B-|\mathcal{C}_{k}|$; $\triangleright$ Update remaining size
@@
18:function DynamicSelect($B,k$):
19:  if $|\mathcal{D}_{k}|\leq\lfloor B/{k}\rfloor$ then:
@@
22:   $n_{k}\leftarrow\lfloor B/{k}\rfloor$;
@@
A model trained on $\mathcal{C}$ is expected to achieve trajectory prediction performance comparable to that obtained from training on the full dataset $\mathcal{D}$. Our method is formally described in Algorithm 1. Next, we explain the key algorithmic components in our methods.
Data Partitioning. As pointed out in Section 1, data imbalance in agent density is a common characteristic of datasets in the autonomous navigation domain. The impacts of such an imbalance are particularly pronounced when Transformer architectures or GNNs that rely on self-attention or graphs to capture dependencies among agents (Zhou et al., 2022c; 2023; Tang et al., 2024; Zhang et al., 2025). Without density-based partitioning, gradient updates are dominated by abundant low-density scenarios, leading to overfitting on sparse interactions and systematic under-training of high-density scenarios. This imbalance prevents the model from learning robust attention patterns. According to this observation, we first compute a density level $\rho(S_{j})$ for each sample $S_{j}\in\mathcal{D}$ based on the number of agents present. Then, we use a fixed interval $\tau$ to partition the dataset into $K$ disjoint subsets $\mathcal{D}_{k}$ for $k\in[K]$ based on these density values such that $S\in\mathcal{D}_{k}$ if $\rho(S)\in[\rho_{\min}+(k-1)\tau,\;\rho_{\min}+k\tau)$, where $\rho_{min}$ is the minimum density level in $\mathcal{D}$.
@@
where $\mathcal{L}_{reg}$ measures the L2 norm between the best mode prediction and the ground truth trajectory, and $\mathcal{L}_{cls}$ aligns the mode predicted probabilities $\pi$ with the best mode. By backpropagating loss $\mathcal{L}$, we obtain the gradient with respect to $\hat{\mathbf{Y}}$ as follows:
@@
Sample Selection. Recall that $B$ is the target size of the constructed training subset. To ensure fair representation across scene complexities, we adopt a dynamic allocation strategy. For each partition $\mathcal{D}_{k}$, a local budget $n_{k}$ is assigned by the DynamicSelect function (see Algorithm 1). This mechanism prioritizes high-density subsets that naturally contain fewer samples, preventing them from being underrepresented when the global budget is small. Formally, the allocation satisfies:
$\sum_{k=1}^{K}n_{k}=B,\quad n_{k}\geq 0.$ (4)
Given total $K$ disjoint subsets and the corresponding gradient feature vectors $\mathcal{G}_{k}$, we present a submodular gain function $P(\cdot)$ to evaluate the contribution of each sample:
@@
where we use a cosine similarity kernel to measure the similarity between the sample $S_{j}$ and other samples. We then apply a greedy optimization strategy, iteratively selecting the sample as follows:
@@
At each iteration, the selected sample $S^{*}$ is added to $\mathcal{C}_{k}$. This process continues until the number of selected samples reaches the budget $n_{k}$ for subset $\mathcal{D}_{k}$. Greedy optimization guarantees that the selected set $\mathcal{C}$ satisfies:
$P(\mathcal{C})\;\;\geq\;\;\left(1-\tfrac{1}{e}\right)P(\mathcal{C}^{*}),$ (7)
where $\mathcal{C}^{*}$ denotes the optimal subset under target size $B$. Notably, we process the subsets starting with those having higher density levels, as these subsets tend to be scarce and underrepresented. This prioritization is consistent with our dynamic allocation strategy, which ensures that complex, high-density scenarios are not discarded in early pruning and that long-tailed scenarios remain adequately represented. For subsets $\mathcal{D}_{k}$ where $|\mathcal{D}_{k}|\leq n_{k}$, we directly set $\mathcal{C}_{k}=\mathcal{D}_{k}$. Finally, we can yield the target dataset $\mathcal{C}=\bigcup_{k=1}^{K}\mathcal{C}_{k}$. By incorporating gradient-based similarity into submodular greedy selection, our method maximizes coverage of the gradient space while maintaining diversity, producing a smaller dataset that remains representative and informative. In addition, we account for efficiency: the additional cost of sample selection is dominated by gradient computation and submodular updates. The overall computational complexity can be approximated as $\mathcal{O}(\text{selection})=\mathcal{O}(|\mathcal{D}|\cdot d)+\mathcal{O}(B\cdot d)$, where $d$ is the dimension of the gradient feature vectors.
Why Naive Strategies Fall Short. To motivate the necessity of density-aware, gradient-guided selection, we examine several alternative strategies: Re-weighting: assigns larger loss weights to high-density samples while retaining the full dataset; Augmenting: duplicates high-density samples to artificially increase their proportion; High-density+Random preserves all high-density samples first and fills the remainder via random sampling; Epoch-wise: dynamically re-selects half the data at the start of each training epoch. These alternatives highlight critical limitations. Both random down-sampling and epoch-wise re-selection significantly reduce training data, but the former discards informative scenarios (Table 1, line 4) while the latter introduces instability despite added diversity (line 3). Simple re-weighting or duplication naively equalizes sample presence but often introduces redundancy, showing that effective density balancing requires more than adjusting counts (line 1–2).
[…]
Augmenting 220k 0.718 1.106 0.115
Weighting 190k 0.715 1.108 0.114
Epoch-wise 95k 0.752 1.189 0.130
Random 95k 0.741 1.164 0.125
High-density+Random 95k 0.724 1.111 0.117
SSTP 95k 0.704 1.073 0.111
Table 1: Comparison with other selection strategies on Argoverse 1. All methods are evaluated at 50% budget (95k samples) unless otherwise specified.
Preservation of high-density scenarios is intuitively beneficial, yet without a principled mechanism to regulate redundancy and ensure representativeness, gains remain limited (line 5). Importantly, in trajectory prediction, the scarcity of high-density scenarios is critical, as reliable performance requires learning complex multi-agent interactions that cannot be captured by oversampling or naive mixing. A central observation is that competence in complex, high-density interactions transfers naturally to simpler scenarios as shown in Figure 2, whereas the reverse does not hold. This asymmetry motivates SSTP, which partitions by density and uses gradient-based selection to preserve scarce high-density scenarios while avoiding redundant low-density ones.